A company that manufactures automatic blowdown control valves (for applications where boilers are operated unsupervised for 24 to 36 hours) has fixed cost of $260,000 per year and variable cost of $725 per valve. The company expects to sell 12,000 valves per year.

Determine the selling price in order for the company to break even.

The selling price for the company to break even is determined to be $_____ per unit.
Answer #1

Break even price is that for which there is no profit or loss.
So, Profit = Total Revenue - Total cost = 0

Total Revenue = Price*Quantity = 12,000P
Total cost = 260,000 + (725*12,000) = 260,000 + 8,700,000 = 8,960,000

So, 12,000P - 8,960,000 = 0
So, 12,000P = 8,960,000
So, P = 8,960,000/12,000 = 746.67

The selling price for the company to break even is determined to be $__746.67___ per unit.